OpenStudy (anonymous):

Find the length of the diameter of a sphere with a surface area of 804.25. A. 12 B. 4 C. 8 D. 16

OpenStudy (amistre64):

do we have a formula for surface area?

OpenStudy (anonymous):

4 pi r^2

OpenStudy (amistre64):

good, then if we rearrange that to get the "r" all by itself then we can determine the diameter of the sphere; since diam is twice r

OpenStudy (amistre64):

Sa = 4pi r^2 Sa/4pi = r^2 sqrt(Sa/4pi) = r 2*sqrt(Sa/4pi) = 2*r = diameter

OpenStudy (amistre64):

i believe thats about the sqrt of 225; so its a little bigger than 15

OpenStudy (anonymous):

one question is a cone a sphere?

OpenStudy (amistre64):

a cone is a cone; and a sphere is a sphere.

OpenStudy (anonymous):

thanks